1. Solve the problem.
The function Y(x) = 48.94 ln can be used to estimate the number of years Y(x) after 1980 required for a certain country's population to reach x million people. In what year will the country's population reach 11 million?
2. Given that log 2 ≈ 0.301 and log 3 ≈ 0.477, find the following.
log6 27
3. Solve. Where appropriate, include approximations to the nearest thousandth. If no solution exists, state this.
ln x - ln (x - 4) = ln 4
4. Solve. Where appropriate, include approximations to the nearest thousandth. If no solution exists, state this.
e^t = 1000

Answer by edjones(8007)

3)
ln x - ln (x - 4) = ln 4
e^(ln (x )) /e^ (ln (x - 4)) =e^( ln (4))
x/(x-4)=4
x=4(x-4)
x=4x-16
-3x=-16
x=16/3
.
Ed